OpenStudy (anonymous):

Which expression does not result in a value of eighty-eight point four one? fifty-three point nine plus thirty-four point five one ninety point six three minus two point two two three hundred point four minus two hundred eleven point nine nine seventy point one plus eighteen point four

OpenStudy (anonymous):

@Jadeishere

OpenStudy (anonymous):

@texaschic101

OpenStudy (jadeishere):

second one, I think

OpenStudy (texaschic101):

(1) 53.9 + 34.51 = 88.41 (2) 90.63 - 2.22 = 88.41 (3) 300.4 - 211.99 = 88.41 (4) 70.1 + 18.4 = 88.5 <-- this one

OpenStudy (jadeishere):

Oh, I read that wrong, lol